The correction of the peltier modules which cool the crystals of the slave laser were slightly increasing (plot 1 ,saturation when reaching 0).
This must be due to the water flow which is somehow getting blocked into the laser head.
To prevent any saturation of the correction of the peltier, this morning at 10:23 UTC, we went in the EER and switched the chiller from the main one to the spare one (which as a reverse flow and a filter) to try to clean up the tubes and the head.
We let the spare chiller run till 12:37 UTC and switched again to the main one. (plot 2).
It seems that this action brought a general improvment (plot 3).
We will monitor the situation and especially the corrections of the peltier modules in the coming days.